OpenStudy (anonymous):

Given the equation -4x + 10 = 2, which order of operations completely solves for x?

OpenStudy (anonymous):

-4x + 10 = 2 -4x + 10 - 10 = 2 - 10 -4x = - 8 -4x/-4 = -8/-4 x = 2 sorry coding went wrong o.o

OpenStudy (anonymous):

@Shelby1997

OpenStudy (anonymous):

is that right ?

OpenStudy (anonymous):

yep